What is the pH of a 2.4x10-4 M NH3 solution? (Kb = 1.8x10-5 )


1
Expert's answer
2021-05-03T07:25:50-0400

NH3 reacts with water to form NH4OH

NH4OH is a weak base.

NH4OH(aq) = NH4+(aq) + OH-(aq)

Kb = [NH4+]×[OH-]/[NH4OH]

or,Kb= [OH-]^2/[NH4OH]

or,1.8×10^-5 = [OH-]^2/2.4×10-4

or,[OH-]^2= 4.32×10-9

or,[OH-] = 6.57×10-5

pOH = -log[OH-]

= -log(6.57×10-5)

= 4.18

Therefore, pH= 14–4.18

= 9.82

Hence the PH of NH3 solution is 9.82